OpenStudy (mkapad01):

Rewrite each trig expression in a simpler form using trig identities. a. 2sin5pcos5p b. sin(pi/6)cos(pi/4)-cos(pi/6)sin(pi/4).

Directrix (directrix):

a. 2sin5pcos5p --- Try the double angle formula for sine on this one.

Directrix (directrix):

sin(pi/6)cos(pi/4)-cos(pi/6)sin(pi/4) Apply the difference formula for sine on this one. http://home.windstream.net/okrebs/page101.html

OpenStudy (mkapad01):

Thanks, but I don't know how to apply these formulas to these problems

Directrix (directrix):

The person I was helping left me so let's see about this.

Directrix (directrix):

What is the double angle formula for sine?

Directrix (directrix):

sin (2x) = 2* sin x * cos x = 2* sin 5p * cos 5p In your problem, x = 5p

Directrix (directrix):

So, sine of what angle will equal the expression you were given? @mkapad01

Directrix (directrix):

If you don't respond, I cannot continue to help.

OpenStudy (mkapad01):

sorry, I had to leave

Directrix (directrix):

We were looking at sine of 2x.

Directrix (directrix):

sin (2x) = 2* sin x * cos x = 2* sin 5p * cos 5p So, sin( 2 * 5p) = 2* sin 5p * cos 5p What is 2 * 5p ?

OpenStudy (mkapad01):

I'm not sure